OpenStudy (anonymous):

z varies jointly as x and y, and inversely as w. If z = 3 when x = 3, y = -2, and w = 4, find z when x = 6, y = 7, and w = -4.

OpenStudy (anonymous):

21?

hartnn (hartnn):

z varies with x z = kx z varies with y z= ky z varies inversely with w z = k/w combining all these z = kxy/w got this ?

OpenStudy (anonymous):

Is the answer 21?

hartnn (hartnn):

z = kxy/w and z = 3 when x = 3, y = -2, and w = 4 from these , can you try to get the value of k ?

OpenStudy (anonymous):

to my knowledge yes. but ask hartnn as well

hartnn (hartnn):

k comes out to be -2

hartnn (hartnn):

and yes, 21 its is :)

OpenStudy (anonymous):

^hartnn why is it to be the same k for all the three x,y,w

OpenStudy (anonymous):

Okay thanks!!!

hartnn (hartnn):

because there will only be one equation and so only one constant is needed. you can take 3 different constants, but when combining , you will combine those 3 constants to one constant.

OpenStudy (anonymous):

^yes you are right hartnn. But it should not necessarily be the same k

hartnn (hartnn):

yes, correct. i just diddn't want to complicate things...

OpenStudy (anonymous):

True that it gives final single k, but having the same k for first 3 equations can mislead someone just trying to figure out the math out of it. Just saying

hartnn (hartnn):

true. sorry.